OpenStudy (nathanjhw):

suppose f(x)=3x^4-2x^2+3x. Then f''(x) =

OpenStudy (solomonzelman):

can find f'(x) first?

OpenStudy (solomonzelman):

(use the power rule)

OpenStudy (anonymous):

??????????????????????????????

OpenStudy (nathanjhw):

@freckles

OpenStudy (solomonzelman):

apply the power rule to each of the terms, 1) what is the derivative of \(\large\color{black}{ 3x^4}\) ? 2) what is the derivative of \(\large\color{black}{ 2x^2}\) ? 3) what is the derivative of \(\large\color{black}{ 3x}\) ?

OpenStudy (solomonzelman):

after you find the f '(x), differentiate f'(x) (also by applying the power rule to each of the terms) and there will be your f''(x).

OpenStudy (solomonzelman):

The last thing you need is, POWER RULE: \(\LARGE\color{black}{ \frac{d}{dx}~x^n=n~x^{n-1}}\) (where n is a constant (but not 0) , and x is the variable) when you have a constant c times the x^n, you get, \(\LARGE\color{black}{ \frac{d}{dx}~cx^n=cn~x^{n-1}}\)

OpenStudy (anonymous):

ohhhhhhhhhhhhhhh so its ^^^^^^^^^^^^^?

OpenStudy (nathanjhw):

I got 12x^3-4x+3x as my answer, but I think I might have done something wrong.

OpenStudy (solomonzelman):

close.

OpenStudy (solomonzelman):

the derivative of 3x is just 3, not 3x.

OpenStudy (nathanjhw):

Oh I meant 12^3- 4x +3

OpenStudy (nathanjhw):

12x^3 -4x +3

OpenStudy (solomonzelman):

YES. So, when \(\Large\color{red}{ f(x)=3x^4-2x^2+3x}\) then \(\Large\color{red}{ f~'(x)=12x^3-4x+3}\)

OpenStudy (solomonzelman):

So after you have, \(\Large\color{red}{ f~'(x)=12x^3-4x+3}\) find f ''(x) by (again) applying the power rule to each term (one by one)....

OpenStudy (anonymous):

yes solomon

OpenStudy (solomonzelman):

1) the derivative of \(\Large\color{blue}{ 12x^3}\) is ? 2) the derivative of \(\Large\color{blue}{ -4x}\) is ? 3) the derivative of \(\Large\color{blue}{ 3}\) is ?

OpenStudy (anonymous):

3